OpenStudy (anonymous):

Solve the system of linear equations below. 2x + y = 5 x + y = 4 answers: x = 1, y = 3 x = 1, y = 5 x = 3, y = 1 x = 9, y = -5

Parth (parthkohli):

Subtract second equation from first.

Parth (parthkohli):

\[x = 5 - 4 = 1\]

Parth (parthkohli):

Now you plug in the value of \(\rm x\) into the second equation.\[\rm 1 + y = 4\]

Parth (parthkohli):

What do you get as a solution to the above?

OpenStudy (anonymous):

x=1, y=4?

Parth (parthkohli):

Erm... nope

Parth (parthkohli):

What is the solution to \(\rm y + 1 = 4\)?

OpenStudy (anonymous):

idk

Parth (parthkohli):

What plus 1 equals 4?

OpenStudy (anonymous):

3

Parth (parthkohli):

Yes, so \(\rm y = 3\).